Journalist Report – May 5th

Crew 235 Journalist Report 05May2021
Authors: Thomas Quayle and Randall Gibson, Crew Journalist

Sol 3: Generally Speaking, It Goes Like This

Morning began with a brisk wind and another species of Rodentia captured. The crew has been remarkably passive about the recurring nature of the event. What began as a mixed curiosity has evolved into a morning routine, complete with EVA procedures, and smiles.

However, the first order of the day: coffee and another round of lively discussion surrounded the breakfast table. Today’s topic was the politics of Earth. We are hopeful they will come to solve their problems in time. The morning fair consisted of sausage and eggs. It has been great to enjoy the variety of meals that each of the crew members has been willing to jump in and prepare.

Little did the crew suspect that the quiet morning would not remain so for long. Even the most well-planned EVA cannot account for everything a crew can prepare for or expect. Their mission was to search for specific geologic samples in a region that was located nearly twice as far as the previous day’s activities.

More than an hour had passed from when the first EVA team had departed when a call came over the radio, “Hab, Hab!..”. The other half of the crew at the Hab quickly responded and was informed that the EVA was being aborted. Within minutes of the radio communique, the EVA team had arrived.

Thankfully no one had been injured, but a malfunction in the power pack of the Commander’s spacesuit (redacted) had occurred that seriously affected the ability for air to be recirculated. While it is not yet known where the malfunction took place or how it was missed, the event has impacted our EVA assessments for the duration of our mission. Each team will now have to continue with one crew member volunteering to abstain from participation. In the end, the second EVA team was able to successfully manage the completion of objectives scheduled for the day.

After the second EVA returned to the Hab, the crew reconvened as a single group to rest and clean the remains of lunch until the next event of the day. Shannon led a crew discussion on the assembly of Estes Rockets and the pitfalls to consider when introducing the topic to students. Each member of the crew was able to construct their own prototype, which should be available for launch tomorrow.

The crew then joined Shannon in the Science Dome to compare the samples that had been collected across multiple EVA’s. In the last significant event of the day, the astronomy box was retrieved from the outpost. This was the final component required to activate the solar observatory, which the crew is looking forward to using tomorrow.

As a social component, the crew finalized a Lego build that had been started the previous day. The genders and position of the Lego figures were debated and rearranged by the crew. The dinner discussion covered several topics, including the similarities to the Lewis and Clark expedition.

Jennifer prepared an incredible culinary masterpiece of southern fried chicken, fried potatoes, biscuits, and gravy by using dehydrated ingredients. This was surely the best meal yet to be prepared on this planet, and she shall be henceforth be known as the “Mars”tha Stewart.

Sol Summary – May 5th

Crew 235 Sol Summary Report 05-May-2021

Sol: 3

Summary Title: Rocks and Rockets!

Author’s name: Atila Meszaros

Mission Status: Crew Biologist

Sol Activity Summary:

We continued or geology hunting as EVA team #1 and #2 went South and reached the 1101 road to collect granite, lava in its different color variation, asphalt, devil’s toenail, and more to be analyzed by our expert, Dr. Shannon Rupert.

The first EVA encountered a small incident (check the EVA report for full explanation), but we are very close to finishing our geology hunt and gathering.

After lunch, we decided to start building our rockets and launching system for tomorrow’s project! Jennifer and Krysta are making dinner as we prepare the reports and get ready for a great night trying new activities and games.

Look Ahead Plan: Rock hunting EVAs in the morning, rocket launching, and going out of sim!

Anomalies in work: None.

Weather: Fully sunny, but with a great temperature throughout the day.

Crew Physical Status: Healthy. Finally learned how to work with dehydrated food and happy with the great meals and bread.

EVA: Tomorrow EVAs #5 and #6

Reports to be filed: Operations Report, EVA Report, EVA Request, Crew Journalist

Support Requested: Nothing to report.

Operations Report – May 5th

Crew 235 Operations Report 5-MAY-2021

SOL: 3

Name of person filing report: Atila Meszaros

Non-nominal systems: 1 EVA suit, Science Dome door, and inner airlock door of the RAM

Notes on non-nominal systems: Shannon used a screwdriver, a drill, a socket set, and a sledgehammer to fix both doors!

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 98% at 18:23 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 71% 2. 80%

Water (loft tank): 50 gallons

Water Meter: 150994.0 units

Water (static tank): 350 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 209.2

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#3 and EVA#4

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0

Notes on ATVs: ATVs will be filled with ethanol-free gas to be ready for tomorrow’s EVA.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #3.

Summary of GreenHab operations: Door not quite fixed.

Summary of ScienceDome operations: Door fixed.

Summary of RAM operations: Door fixed.

Summary of any observatory issues: The rotation problem was a battery issue.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight!

EVA Report – May 5th

Crew 235 EVA Report 05 May 2021

EVA # 4, 5

Author: Jen Carver-Hunter

Purpose of EVA: Continuation of rock sample collection

EVA 4 Start time: 0900

EVA 4 End time: 1000

EVA 5 Start Time: 1200

EVA 5 End time: 1430

Narrative: One of our crew’s goals is to collect rock samples that replicate the rocks found on Mars. For today’s EVA, our targeted specimens were lava, basalt, and granite. In addition, we developed a scavenger hunt challenge for the two EVA crews to use as a competition that included specimens of gryphaea, purple lava, hematite, and sandstone blueberries.

The first planned stop for the EVA was the route to Barainca Butte, and the second planned stop on return to the Hab was east of Zubrin’s head.

The first EVA crew left the hab in vehicles at 0900 and followed Cow Dung Road south to Road 1101. After following Road 1101 west toward Barainca Butte, the crew stopped to collect rock samples, which included lava, granite, and basalt, and gryphaea. The Commander’s suit battery discharged so the EVA was aborted and the crew returned to the hab at 1000.

Because that crew’s EVA mission was not completed they requested that the second EVA crew collect extra samples to meet the day’s goals. They also requested that the second crew return some unwanted petrified wood samples to the west side of Kissing Camel Ridge, as was suggested by Dr. Rupert when she approved today’s EVA last night. Due to the dead battery in Spacesuit 7, the second EVA was only able to accommodate 3 crew members. Krysta volunteered to stay at the hab and requested that her crewmates bring samples back for her collection.

The second EVA crew left the hab in vehicles at 1200 and also followed Cow Dung Road south to Road 1101. They followed Road 1101 toward Barainca Butte and collected their samples along the road. After collecting on Road 1101, the EVA team turned around and turned north on Cow Dung Road until they reached “the Squirrel” landmark. They followed a wash to the west (they were east of Zubrin’s head) until they found the sandstone blueberry nursery.

At this point, they were able to complete the second objective of their mission. The final objective was to return the pieces of petrified wood they were done using to their collection site. Crew stopped south of Kissing Camels Ridge and walked west to return their samples before returning back to the Hab. The second EVA crew returned to the Hab with no difficulty and shared their collected samples with the first EVA crew.

Destination: Barainca Butte, east of Zubrin’s head, Kissing Camel Ridge

Coordinates (use UTM NAD27 CONUS): Destination 1 – 519700, 4247200; Destination 2 – 519750, 4248000; Destination 3 – 518250, 4249250

EVA 4 Participants: Jen, Krysta, Jennifer, Jeff

EVA 5 Participants: Atila, Randall, Thomas

Road(s) and routes per MDRS Map: Entrance Road to Cow Dung Road to Road 1101

Mode of travel: Pathfinder Rover and 2 ATVs.

Copyright © The Mars Society. All rights reserved. | Main Site